OpenStudy (mathhelp346):

Graph the function y=x(125-x). Find the dimensions that maximize the area of the pen. (What are the coordinates to the highest point in the graph?)

OpenStudy (callisto):

y=x(125-x) = -x^2 + 125x = -(x^2 -125x + 62.5^2) -(-62.5^2) = -(x-62.5)^2 + 3906.25 So, vertex is at (62.5, 3906.25)

OpenStudy (mathhelp346):

how did you get this line = -(x^2 -125x + 62.5^2) -(-62.5^2)

OpenStudy (anonymous):

is y a model for the area of a pen?

OpenStudy (mathhelp346):

yes